Exercise 4-12 Equivalent Units; Assigning Costs; Cost Reconciliation-Weighted Average Method (L04-2, LO4-4, LO4-5) Superior MPlease answer all parts! Thanks!
Answer #1

1) Equivalent unit

Material Labor Overhead
Equivalent unit of ending WIP 3400*50% = 1700 3400*30% = 1020 1020

2) Cost of ending work in process

Material Labor Overhead Total
Cost of ending work in process 1700*14.40 = 24480 1020*3.8 = 3876 1020*8.20 = 8364 36720

3) Cost of unit transferred out

Material Labor Overhead Total
Cost of Unit transferred out 25600*14.40 = 368640 25600*3.8 = 97280 25600*8.20 = 209920 675840

4) Cost reconciliation

Total cost to be accounted for 712560
Cost accounted as
Cost of unit transferred out 675840
Cost of ending WIP 36720
Total cost accounted for 712560